Make x the subject of the following:
y = 2 - x² / 2x² + 3
y = √2x- 3y (the sign before 2x is of square root)
1/a + 1/√x = y

Answer by stanbon(75887)

Make x the subject of the following:
y = 2 - x² / 2x² + 3
y(2x^2+3) = 2-x^2
2x^2y+x^2 = 2-3y
x^2(2y+1) = 2-3y
x^2 = (2-3y)/(2y+1)
-------------------------------------
y = √2x- 3y (the sign before 2x is of square root)
sqrt(2x) = y + 3y
2x = (4y)^2
x = 8y^2
------------------------------
1/a + 1/√x = y
1/sqrt(x) = y-(1/a)
1/sqrt(x) = (ay-1)/a
sqrt(x) = a/(ay-1)
x = [a/(ay-1)]^2
====================================
Cheers,
Stan H.